# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ 4 Examples of loss functions
Now, we instantiate Theorem 5 for several examples of Fenchel–Young losses to discuss the convergence rate. Instead of specifying a loss function $\ell(z)=\phi^{*}(-z)$, we directly specify its potential function $\phi$ subsequently. For each $\phi$, we compute $(\alpha,C_{\phi})$ in (3) to investigate the convergence rate given by Theorem 5, by leading $\bar{\varepsilon}$ (and thus $\bar{\mu}$) vanishingly small. In addition, we can compute separation margin $m$ by Proposition 4 if exists; otherwise, we need to compute $\rho$ for a loss (see Lemma 23). Table 1 summarizes different loss functions and their GD convergence rates. All the detailed calculations are deferred to Appendix E, where we have an additional example of $\phi$ (pseudo-spherical entropy) with non-converging $\alpha$.
@@
Shannon $\infty$ $1/4$ $1$ $\widetilde{\Omega}(\varepsilon^{-1})$
Semi-circle $\infty$ $1/4$ $2$ $\Omega(\varepsilon^{-4})$
Tsallis $(0,1)$infty $\dfrac{2^{q-3}}{q}$ $\dfrac{1}{q}$ $\Omega(\varepsilon^{-2/q})$
@@
$(2,\infty)$infty $1/2$ $\Omega(\varepsilon^{-1/2})$
Rényi $(0,1)$infty $1/4q$ $\dfrac{1}{q}$ $\Omega(\varepsilon^{-2/q})$
@@
$2$infty $1/3$ $\Omega(\varepsilon^{-1/3})$
@@
Consider the binary Shannon (neg)entropy $\phi(\mu)=\mu\ln\mu+(1-\mu)\ln(1-\mu)$. The generated Fenchel–Young loss is the logistic loss $\ell(z)=\ln(1+\exp(-z))$, which enjoys the self-bounding property and hence does not have separation margin (see Appendix D). The loss parameters are $\alpha=1$ and $C_{\phi}=1$. Moreover, we know $C_{g}=1$ and $\rho(\lambda)\leq 1+\ln^{2}(\lambda)$ [55]. Plugging this back to Theorem 5, we have the $\varepsilon$-optimal risk at most after
$T\gtrsim\left[\frac{4\sqrt{2}(\log_{2}(\gamma^{2}\eta)+1)}{\eta}+\frac{1}{\ln 2}\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}}\quad\text{iterations,}$,
where logarithmic factors in $\varepsilon^{-1}$ are ignored. This indicates the rate $T=\widetilde{\Omega}(\varepsilon^{-1})$, recovering the standard GD convergence rate under the stable regime but with arbitrary stepsize $\eta$. In Section 5, we compare this rate with Wu et al. [55] in more detail.
@@
Consider $\phi(\mu)=-2\sqrt{\mu(1-\mu)}$. The generated Fenchel–Young loss (we call the semi-circle loss) $\ell(z)=(-z+\sqrt{z^{2}+4})/{2}$ enjoys the self-bounding property and does not have separation margin since $\phi^{\prime}(\mu)\to-\infty$ as $\mu\downarrow 0$ (see Appendix D). The semi-circle loss is relevant to the exponential/boosting loss $\ell_{\text{exp}}(z)=\exp(-z)$, which has the semi-circle entropy as the Bayes risk [13, 1]. The loss parameters are $\alpha=2$ and $C_{\phi}=1$. Moreover, we have $C_{g}=1$ and $\rho(\lambda)\leq 5\lambda/(2\ln\lambda)$. Plugging this back to Theorem 5, we have the $\varepsilon$-optimal risk at most after
$T>\underbrace{\frac{40n^{6}}{\gamma^{2}\eta\ln(2\gamma^{2}\eta)}\varepsilon^{-4}}_{\text{extra price for lacking separation margin}}+\frac{2n^{3}}{\gamma^{2}}\varepsilon^{-2}\quad\text{iterations,}$,
@@
This convergence rate of the semi-circle loss is even worse than the GD convergence rate for general convex smooth functions, $T=\Omega(\varepsilon^{-1})$. This is because the perceptron argument is merely sufficient for GD convergence. Nonetheless, the perceptron argument more informatively states that we have $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle/\|\mathbf{w}_{t}\|\gtrsim\varepsilon^{\alpha}$ after minimizing the risk at the $\varepsilon$-optimal level—by combining the inequalities (,)($\clubsuit$,$\diamondsuit$)( ♣ , ♢ ) (in Eq. 4). This indicates that the loss function with larger $\alpha$ yields slower parameter alignment toward $\mathbf{w}_{*}$.
@@
For $q>0$ with $q\neq 1$, consider the Tsallis $q$-(neg)entropy
@@
  • When $10<q<$ 1: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ does not have separation margin.
  • When $1<q\leq 2$: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ has separation margin $m=q/(q-1)$.
  • When $2<q$: $(\alpha,C_{\phi})=(1/2,\sqrt{2/q})$, and $\phi^{*}$ has separation margin $m=q/(q-1)$.
For all cases, $\alpha$ and $C_{\phi}$ stay in $(0,\infty)$. The convergence rate is $T=\Omega(\varepsilon^{-2/q})$ for $q\in(0,1)$ (by Corollary 24); $T=\Omega(\varepsilon^{-1/q})$ for $q\in(1,2)$; $T=\Omega(\varepsilon^{-1/2})$ for $2\leq q$. This suggests that we have a better convergence rate over the Shannon case when $q>1$ and the best rate is $\Omega(\varepsilon^{-1/2})$.
@@
For $q\in(0,2]\setminus\left\{{1}\right\}$, consider the Rényi $q$-(neg)entropy
@@
  • When $10<q<$ 1: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ does not have separation margin.
  • When $21<q<$ 2: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ has separation margin $m=q/(q-1)$.
@@
For all cases, $\alpha$ and $C_{\phi}$ stay in $(0,\infty)$. The convergence rate is $T=\Omega(\varepsilon^{-2/q})$ for $q\in(0,1)$ (by Corollary 24); $T=\Omega(\varepsilon^{-1/q})$ for $q\in(1,2)$; $T=\Omega(\varepsilon^{-1/3})$ for $q=2$. Surprisingly, we have a “leap” of the order from $\varepsilon^{-1/q}$ to $\varepsilon^{-1/3}$ as $q\uparrow 2$, and the convergence rate $\Omega(\varepsilon^{-1/3})$ is far better than the Shannon and Tsallis cases. When $q=2$, Corollary 6 implies that we have the $\varepsilon$-optimal risk at most after
$T>\sqrt[3]{8/3}\frac{n^{4/3}}{\gamma^{2}}\left(\frac{8}{\eta}+1\right)\varepsilon^{-1/3}\quad\text{iterations.}$.
@@ 5 Discussion and open problems
Phase transition.
Assumption 4.
  1. A.
  2. B.
Theorem 7 (Wu et al. [55]).
Consider (GD) with stepsize $\eta>0$ and initialization $\mathbf{w}_{0}=\mathbf{0}$ under a loss $\ell$ satisfying Assumptions 3B and 3C, and 4A. Let $T$ be the maximum number of steps. Then, we have the following:
  • The EoS phase. For every $t>0$, we have
  • The stable phase. If $s<T$ is such that
    If $\tau\leq T$, (6) holds for some $s\leq\tau$. Moreover, if $\ell$ additionally satisfies Assumption 4B and $\eta\geq 1$, there exists $C_{2}>0$ that depends on $C_{e}$, $C_{g}$, $C_{\beta}$, $\ell(0)$, and $n$ such that $\tau$ is improved as follows:
    $\tau\coloneqq\frac{C_{2}}{\gamma^{2}}\max\left\{{\eta,n}\right\}.$
The proof consists of Lemma 15 (the EoS phase), Lemma 19 (the stable phase), and Lemmas 20 and 21 (phase transition time), respectively, in Appendix C. We restate these proofs for linear models.
@@
To get insight from Theorem 7, consider $\ell$ is the logistic loss. Since the logistic loss satisfies both Assumptions 4A and 4B with $\rho(\lambda)=\widetilde{\Theta}(1)$ [55, Proposition 5], the EoS phase terminates within $\tau=\mathcal{O}(\eta)$ steps at most and converges in the rate $L(\mathbf{w}_{t})=\widetilde{\mathcal{O}}(1/(\eta t))$ under the stable phase. That said, stepsize $\eta$ trades off the phase transition time for the stable convergence rate; if we know the maximum number of steps $T$ in advance, the choice $\eta=\Theta(T)$ balances them, achieving the acceleration to $L(\mathbf{w}_{t})=\widetilde{\mathcal{O}}(1/T^{2})$. This is arguably a groundbreaking instance to demonstrate how GD benefits from large stepsize and why EoS is necessary therein.
@@
Nevertheless, we would like to highlight two caveats. First, we must ensure $L(\mathbf{w}_{s})\leq\ell(0)/n$ in (6) before entering the stable phase. This means that our linear model has already classified all points correctly during the EoS phase since any single point $\mathbf{z}_{i}$ incurs loss at most $\ell(\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle)\leq\ell(0)\implies\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ (cf. Lemma 18 in Appendix C).222 Cai et al. [14] extends Wu et al. [54] for two-layer near-homogeneous NNs, where it is not explicit that the model correctly classifies all points after the EoS phase. Taking a closer look, we can see that their Lemma A.7 leverages the well-controlled risk, which is an alternative expression to “$L(\mathbf{w}_{s})\leq\ell(0)/n$” under their setup. GD keeps improving the logistic loss after the stable phase just because the logistic loss does not enjoy separation margin and never touches strict zero. We refer interested readers to the relevant discussion in Tyurin [50], who argues that the stable loss improvement might be due to the artifact of the logistic loss.
@@
Second, (6) suggests that the risk must be once $\mathcal{O}(1/T)$-optimal (with the optimally balancing choice $\eta=\Theta(T)$) before benefitting from the super-fast rate $\widetilde{\mathcal{O}}(1/T^{2})$. Yet, GD under some loss functions including the Tsallis $q$-loss ($q>1$) and the Rényi $q$-loss ($q>1$) achieves better risk with the same GD steps. If our goal is simply to classify all training points, these alternative losses might do better jobs in terms of optimization solely.
@@
Figure 2: Under the same setup as Figure 1, we show $\|\mathbf{w}_{t}\|$ along the number of steps $t$ with different loss functions.
@@
Having said that, the phase transition may play an important role in implicit bias. Wu et al. [54] shows under the linearly separable case that logistic regression optimized with large-stepsize GD enlarges the norm $\|\mathbf{w}_{t}\|$ toward the max-margin direction in rate $\Omega(\ln(t))$. Thus, we may argue that $\mathbf{w}_{t}$ gradually comes to classify all data points correctly during the EoS phase and evolves toward the max-margin direction in the stable phase. Such a max-margin implicit bias is ubiquitous for separable logistic regression [47], and is demonstrated even for a nonlinear model as well [14].
We reported how $\|\mathbf{w}_{t}\|$ evolves under the pilot setup in Figure 2 with different loss functions. As seen, the logistic loss inflates $\|\mathbf{w}_{t}\|$ endlessly as well as the Tsallis $0.5$-loss, which does not have separation margin and satisfies the self-bounding property. In stark contrast, the Tsallis $2$-loss prevents $\|\mathbf{w}_{t}\|$ from growing endlessly just because of its separation margin—recall the norm upper bounds of the norm $\|\mathbf{w}_{t}\|$ in (12) and the growth rate $\rho$ in Lemma 10. This raises two open questions: (1) Do we have similar implicit bias aligning toward the max-margin direction under a loss function with separation margin? (2) What are benefits and caveats of endless growing of $\|\mathbf{w}_{t}\|$? The latter is interesting in particular because excessively large $\|\mathbf{w}_{t}\|$ may practically lead to overconfidence issue [53] and worse generalization due to prohibitively large within-class variance [21].
The study on implicit bias beyond the logistic loss (or those with the self-bounding property) has been very scarce. To our knowledge, Lizama [30] crafted the complete hinge loss, which behaves like the hinge loss before GD converges to the zero risk yet incurs an extra penalty to artificially align the parameter toward the max-margin direction—Lizama [30, Figure 2.1] illustrates this idea clearly. Together with the benefits and caveats of the max-margin implicit bias, we believe this is an interesting open topic.
@@
Last but not least, we may potentially have another benefit of loss functions with separation margin. Take a look at Figure 1 again. Loss functions without separation margin, such as the Tsallis $1.5$- and $2$-losses, converge to exact zero within finite time when sufficiently large stepsize is used. Such finite-time convergence under the linearly separable case can be shown without significant challenges if we use perceptron, or even the hinge loss, while becoming highly non-trivial in case of twice-differentiable loss functions. This is because the perceptron argument requires a non-degenerate lower bound on $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ (see (4)), which is not straightforward therein as the loss gradient can be arbitrarily small positive (due to the twice differentiability of the loss). We conjecture that an additional data assumption is necessary because the loss gradient could be adversarially vanishing against GD convergence, and leave this as future work.
Acknowledgment
HB is supported by JST PRESTO (Grant No. JPMJPR24K6). SS is supported by JST ERATO (Grant No. JPMJER1903). YT is supported by JSPS KAKENHI (Grant No. 23KJ1336).

### v2

@@ 4 Examples of loss functions
Now, we instantiate Theorem˜5 for several examples of Fenchel–Young losses to discuss the convergence rate. Instead of specifying a loss function $\ell(z)=\phi^{*}(-z)$, we directly specify its potential function $\phi$ subsequently. For each $\phi$, we compute $(\alpha,C_{\phi})$ in (3) to investigate the convergence rate given by Theorem˜5, by taking $\bar{\varepsilon}$ (and thus $\bar{\mu}$) vanishingly small. In addition, we can compute separation margin $m$ by Proposition˜4 if exists; otherwise, we need to compute $\rho$ for a loss (see Lemma˜27). Table˜1 summarizes different loss functions and their GD convergence rates. All the detailed calculations are deferred to Appendix˜F, where we have an additional example of $\phi$ (pseudo-spherical entropy) with non-converging $\alpha$.
@@
Shannon $\infty$ $1/4$ 11 $\widetilde{\Omega}(\varepsilon^{-1})$
Semi-circle $\infty$ $1/4$ 22 $\Omega(\varepsilon^{-4})$
Tsallis $(0,1)$ $\infty$ $\dfrac{2^{q-3}}{q}$ $\dfrac{1}{q}$ $\Omega(\varepsilon^{-2/q})$
@@
$(2,\infty)$ $\infty$ $1/2$ $\Omega(\varepsilon^{-1/2})$
Rényi $(0,1)$ $\infty$ $1/4q$ $\dfrac{1}{q}$ $\Omega(\varepsilon^{-2/q})$
@@
22 $\infty$ $1/3$ $\Omega(\varepsilon^{-1/3})$
@@
Consider the binary Shannon (neg)entropy $\phi(\mu)=\mu\ln\mu+(1-\mu)\ln(1-\mu)$. The generated Fenchel–Young loss is the logistic loss $\ell(z)=\ln(1+\exp(-z))$, which enjoys the self-bounding property and hence does not have separation margin (see Appendix˜E). The loss parameters are $\alpha=1$ and $C_{\phi}=1$. Moreover, we know $C_{g}=1$ and $\rho(\lambda)\leq 1+\ln^{2}(\lambda)$ [63]. Plugging this back to Theorem˜5, we have the $\varepsilon$-optimal risk at most after
$T\gtrsim\left[\frac{4\sqrt{2}(\log_{2}(\gamma^{2}\eta)+1)}{\eta}+\frac{1}{\ln 2}\right]\frac{n\varepsilon^{-1}}{\gamma^{2}}\quad\text{iterations,}$
where logarithmic factors in $\varepsilon^{-1}$ are ignored. This indicates the rate $T=\widetilde{\Omega}(\varepsilon^{-1})$, recovering the standard GD convergence rate under the stable regime but with arbitrary stepsize $\eta$. In Section˜5, we compare this rate with [63] in more detail.
@@
Consider $\phi(\mu)=-2\sqrt{\mu(1-\mu)}$. The generated Fenchel–Young loss (we call the semi-circle loss) $\ell(z)=(-z+\sqrt{z^{2}+4})/{2}$ enjoys the self-bounding property and does not have separation margin since $\phi^{\prime}(\mu)\to$-$\infty$ as $\mu\downarrow 0$ (see Appendix˜E). The semi-circle loss is relevant to the exponential/boosting loss $\ell_{\text{exp}}(z)=\exp(-z)$, which has the semi-circle entropy as the Bayes risk [15, 1]. The loss parameters are $\alpha=2$ and $C_{\phi}=1$. Moreover, we have $C_{g}=1$ and $\rho(\lambda)\leq 5\lambda/(2\ln\lambda)$. Plugging this back to Theorem˜5, we have the $\varepsilon$-optimal risk at most after
$T>\underbrace{\frac{40n^{6}}{\gamma^{2}\eta\ln(2\gamma^{2}\eta)}\varepsilon^{-4}}_{\text{extra price for lacking separation margin}}+\frac{2n}{\gamma^{2}}\varepsilon^{-2}\quad\text{iterations,}$
@@
This convergence rate of the semi-circle loss is even worse than the GD convergence rate for general convex smooth functions, $T=\Omega(\varepsilon^{-1})$. This is because the perceptron argument is merely sufficient for GD convergence. Nonetheless, the perceptron argument more informatively states that we have $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle/\|\mathbf{w}_{t}\|\gtrsim\varepsilon^{\alpha}$ after minimizing the risk at the $\varepsilon$-optimal level—by combining the inequalities $(\clubsuit,\diamondsuit)$ (in Eq.˜4). This indicates that the loss function with larger $\alpha$ yields slower parameter alignment toward $\mathbf{w}_{*}$.
@@
For $q>0$ with $q\neq 1$, consider the Tsallis qq-(neg)entropy
@@
  • When $0<q<1$: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ does not have separation margin.
  • When $1<q\leq 2$: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ has separation margin m=q/(q1)m=q/(q-1).
  • When $2<q$: $(\alpha,C_{\phi})=(1/2,\sqrt{2/q})$, and $\phi^{*}$ has separation margin m=q/(q1)m=q/(q-1).
For all cases, $\alpha$ and $C_{\phi}$ stay in $(0,\infty)$. The convergence rate is $T=\Omega(\varepsilon^{-2/q})$ for $q\in(0,1)$ (by Corollary˜28); $T=\Omega(\varepsilon^{-1/q})$ for $q\in(1,2)$; $T=\Omega(\varepsilon^{-1/2})$ for $2\leq q$. This suggests that we have a better convergence rate over the Shannon case when $q>1$ and the best rate is $\Omega(\varepsilon^{-1/2})$.
@@
For $q\in(0,2]\setminus\left\{{1}\right\}$, consider the Rényi qq-(neg)entropy
@@
  • When $0<q<1$: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ does not have separation margin.
  • When $1<q<2$: $(\alpha,C_{\phi})=(1/q,1)$, and $\phi^{*}$ has separation margin m=q/(q1)m=q/(q-1).
@@
For all cases, $\alpha$ and $C_{\phi}$ stay in $(0,\infty)$. The convergence rate is $T=\Omega(\varepsilon^{-2/q})$ for $q\in(0,1)$ (by Corollary˜28); $T=\Omega(\varepsilon^{-1/q})$ for $q\in(1,2)$; $T=\Omega(\varepsilon^{-1/3})$ for $q=2$. Surprisingly, we have a “leap” of the order from $\varepsilon^{-1/q}$ to $\varepsilon^{-1/3}$ as $q\uparrow 2$, and the convergence rate $\Omega(\varepsilon^{-1/3})$ is far better than the Shannon and Tsallis cases. When $q=2$, Corollary˜6 implies that we have the $\varepsilon$-optimal risk at most after
$T>\sqrt[3]{8/3}\frac{n}{\gamma^{2}}\left(\frac{8}{\eta}+1\right)\varepsilon^{-1/3}\quad\text{iterations.}$
@@ 5 Discussion and open problems
Comparison with Wu et al. [63].
@@
The large-stepsize logistic regression has been shown to exhibit the following phase transition [63]: the GD sequence initially stays in the EoS phase such that the risk $L(\mathbf{w}_{t})$ fluctuates initially with its average $t^{-1}\sum_{k}L(\mathbf{w}_{k})$ controlled. Once we experience $L(\mathbf{w}_{t})\lesssim\min\left\{{1/\eta,\ell(0)/n}\right\}$, which is possible within $\mathcal{O}(\eta)$ steps at most, GD leaves the EoS and the loss converges in the rate $L(\mathbf{w}_{t})=\widetilde{\mathcal{O}}(1/(\eta t))$. The stepsize $\eta$ trades off the phase transition time for the stable convergence rate. If we know the maximum number of steps $T$ in advance, the choice $\eta=\Theta(T)$ balances them, achieving the acceleration to $L(\mathbf{w}_{t})=\widetilde{\mathcal{O}}(1/T^{2})$. We detail them in Appendix˜D. This is arguably interesting to demonstrate how GD benefits from large stepsize.
@@
Nevertheless, we would like to highlight two caveats. First, we must undergo $L(\mathbf{w}_{t})\leq\ell(0)/n$ before exiting the EoS phase. This means that our linear model has already classified all points correctly during the EoS phase since any single point $\mathbf{z}_{i}$ incurs loss at most $\ell(\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle)\leq\ell(0)\implies\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ (cf. Lemma˜22 in Appendix˜D). [^4] Cai et al. [16] extends Wu et al. [62] for two-layer near-homogeneous NNs, where it is not explicit that the model correctly classifies all points after the EoS phase. Taking a closer look, we can see that their Lemma A.7 leverages the well-controlled risk, which is an alternative expression to $L(\mathbf{w}_{s})\leq\ell(0)/n$” under their setup. GD keeps improving the logistic loss after the stable phase just because the logistic loss does not enjoy separation margin and never touches strict zero. We refer interested readers to the relevant discussion in Tyurin [58], who argues that the faster convergence in the stable phase is attributed to the choice of the logistic loss.
@@
Second, the EoS termination condition $L(\mathbf{w}_{t})\lesssim 1/\eta$ suggests that the risk must be once $\mathcal{O}(1/T)$-optimal (with the optimally balancing choice $\eta=\Theta(T)$) before benefitting from the super-fast rate $\widetilde{\mathcal{O}}(1/T^{2})$. Yet, GD under some loss functions including the Tsallis $q$-loss ($q>1$) and the Rényi $q$-loss ($q>1$) achieves better risk with the same GD steps. If our goal is simply to classify all training points, these alternative losses might do better jobs in terms of optimization solely.
@@
Figure 2: Under the same setup as Figure˜1, we show $\|\mathbf{w}_{t}\|$ along the number of steps $t$ with different losses.
@@
Having said that, the phase transition may play an important role in implicit bias. It was shown under the linearly separable case that logistic regression optimized with GD enlarges the norm $\|\mathbf{w}_{t}\|$ toward the max-margin direction in rate $\Omega(\ln(t))$ [54, 62, 16]. Thus, we may argue that $\mathbf{w}_{t}$ gradually comes to classify all data points correctly during the EoS phase and evolves toward the max-margin direction in the stable phase.
We reported how $\|\mathbf{w}_{t}\|$ evolves under the pilot setup in Figure˜2 with different loss functions. As seen, the logistic and Tsallis $0.5$ losses inflate $\|\mathbf{w}_{t}\|$ endlessly, which do not have separation margin. In stark contrast, the Tsallis $2$-loss prevents $\|\mathbf{w}_{t}\|$ from growing endlessly just because of its separation margin—recall the norm upper bounds of the norm $\|\mathbf{w}_{t}\|$ in (11) and the growth rate $\rho$ in Lemma˜9. This raises two open questions: (1) Do we have similar implicit bias aligning toward the max-margin direction under a loss function with separation margin? (2) What are benefits and caveats of endless growing of $\|\mathbf{w}_{t}\|$? The latter is particularly relevant to the overconfidence issue due to excessively large $\|\mathbf{w}_{t}\|$ [61] and worse generalization due to prohibitively large within-class variance [26]. Wu et al. [64] argues that excessively large $\|\mathbf{w}_{t}\|$ leads to an inconsistent estimator.
The study on implicit bias for loss functions with the self-bounding property has been very scarce. To our knowledge, [35] crafted the complete hinge loss, which behaves like the hinge loss before GD converges to the zero risk yet incurs an extra penalty to artificially align the parameter toward the max-margin direction. Together with the benefits and caveats of the max-margin implicit bias, we believe this is an interesting open topic.
Dependency on $n$.
Our main result (Theorem˜5) provides the rate depending on the factor $n$. This extra factor with respect to $n$ arises due to the worst-case analysis such that we have at least one “bad” direction $\mathbf{z}_{i}$ before the convergence, corresponding to the inequality (9) in the proof of Theorem˜5 (see Appendix˜B). This worst-case scenario supposes that all data points are nearly equidistant, which is unlikely since most data points tend to cluster in similar directions. We conjecture that this $n$-dependency is not essential with additional mild data assumptions.
The stochastic case.
Our result can be extended for the stochastic gradient descent (SGD). Consider the scenario where we sample one fresh data point at each $t$ and update the linear parameter with the loss function computed on this sample. Under the similar setup to Theorem˜5, the population risk is $\varepsilon$-optimal with high probability after $T=\Omega(\varepsilon^{-(\alpha+2)})$. The formal statement and proof are shown in Appendix˜C. While this rate apparently looks significantly slower than the GD rate $T=\Omega(\varepsilon^{-\alpha})$, the extra iterations $\varepsilon^{-2}$ is necessary for collecting sufficient samples to estimate the population risk. By noting that the GD/SGD updates consume nn/one samples, the GD/SGD rates are comparable in terms of the number of consumed samples.
@@
Last but not least, we may potentially have another benefit of loss functions with separation margin. Take a look at Figure˜1 again. Loss functions without separation margin, such as the Tsallis $1.5$- and $2$-losses, converge to exact zero within finite time when sufficiently large stepsize is used. Such finite-time convergence under the linearly separable case can be shown without significant challenges if we use perceptron, or even the hinge loss, while becoming highly non-trivial in the case of twice-differentiable loss functions. This is because the perceptron argument requires a non-degenerate lower bound on $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ (see (4)), which is not straightforward therein as the loss gradient can be arbitrarily small positive (due to the twice differentiability of the loss). We conjecture that an additional data assumption is necessary because the loss gradient could be adversarially vanishing against GD convergence, and leave this as future work.
Acknowledgments
HB is supported by JST PRESTO (Grant No. JPMJPR24K6). SS was supported by JST ERATO (Grant No. JPMJER1903). YT is supported by JSPS KAKENHI (Grant No. 23KJ1336).